A network provider investigates the load of its network. The number of concurrent users is recorded at fifty locations in thousands of connected users.

17.2 22.1 18.5 17.2 18.6 14.8 21.7 15.8 16.3 22.8

24.1 13.3 16.2 17.5 19.0 23.9 14.8 22.2 21.7 20.7

13.5 15.8 13.1 16.1 21.9 23.9 19.3 12.0 19.9 19.4

15.4 16.7 19.5 16.2 16.9 17.1 20.2 13.4 19.8 17.7

19.7 18.7 17.6 15.9 15.2 17.1 15.0 18.8 21.6 11.9
1
Expert's answer
2021-07-12T07:21:44-0400

"\bar{x}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nx_i=17954"

"s^2=\\dfrac{1}{n-1}\\displaystyle\\sum_{i=1}^n(x_i-\\bar{x})^2=\t9968248.9795918"

"s=\sqrt{s^2}=3157.2534"